OpenStudy (anonymous):

an equation for a circle with center (0,0) and radius is 8

OpenStudy (anonymous):

x^2+y^2=8^2

OpenStudy (anonymous):

if centre is (a,b) and radius is r equation is given by... \[(x-a)^{2}+(y-b)^{2}=r^{2}\] try it out with the above values!!1

OpenStudy (anonymous):

well the options are a)x^2+y^2=64 or b)x^2+y^2=8

OpenStudy (anonymous):

you substitute the values in above equation and get answer.......

hero (hero):

Actually if the center is (0,0), then the equation of a circle is x^2 + y^2 = r^2

OpenStudy (anonymous):

soooo would the answer be x^2+y^2= 64 ?

hero (hero):

Yup

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

okay so theres a circle, and the raduis = 7. and it says write an equation, would it be x^2+y^2=7^2 ?

hero (hero):

They didn't give you the coordinates of its center point?

hero (hero):

The equation of a circle is x^2 + y^2 = r^2 if and only if the the coordinates of its center point is (0,0).

OpenStudy (anonymous):

Okay yes, thank you so much ! :)

hero (hero):

But you didn't answer my question

OpenStudy (anonymous):

oh, yeeh the coordinates are (0,0) the center

hero (hero):

Okay, well, then, yes, your equation is correct. You can either say: x^2 + y^2 = 7^2 or x^2 + y^2 = 49

OpenStudy (anonymous):

okay, and theres another circle, and the points are outside the circle, it says (6,1) how would I write that equation ?

hero (hero):

You can use the formula the other guy provided

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=r^2 ?

hero (hero):

Yes, that :)

hero (hero):

Be careful with the negative

OpenStudy (anonymous):

so it would be (0-6)^2+(0-1)^2=r^2 ?

hero (hero):

I had a feeling you were going to do that. You don't replace x with zero by the way

OpenStudy (anonymous):

and when I solve it would be 37=r^2 which would then be r=6.1 ?

OpenStudy (anonymous):

Oh then what do I do ?

hero (hero):

No, you replaced x and y with zeroes which is forbidden.

hero (hero):

The center point is (a,b) not (x,y) in this case. In other words, you leave x and y alone

hero (hero):

Wait a sec...The points are OUTSIDE the circle? You should probably post or draw the picture.

hero (hero):

This is a different case if a point is outside the circle.

OpenStudy (anonymous):

ohh okay hold on

hero (hero):

Usually, the given formulas are used when a point is on the circle and you know the center point.

OpenStudy (anonymous):



OpenStudy (anonymous):

something like that

hero (hero):

And (6,1) is not on the circle?

hero (hero):

Can you figure out what any one of the points I just drew are?